During June, Health Foods manufactures a single product. The Company's material purchases amounted to 9,000 pounds at a price of $9.60 per pound. Actual costs incurred in the actual output production of 3,300 units were as follows:

Direct material:          $67,200 ($9.60 per pound)

Direct labor:                $312,000 ($24.00 per hour)

The standards for one unit of Health Foods Company's product are as follows:

Direct material:

Direct labor:

Quantity, 2 pounds per unit

Quantity, 4 hours per unit

Price, $9.10 per pound

Rate, $22.50 per hour

Required: Showing your calculations, Compute the following variances and indicate whether each variance is favorable or unfavorable:

Requirements & Calculations

Final Answer

1.

Direct-material price variance.

2.

Direct-material quantity variance.

3.

Direct-labor rate variance.

4.

Direct-labor efficiency variance.

please make sure the answer is correct 100%
Answer #1

Required: Showing your calculations, Compute the following variances and indicate whether each variance is favorable or unfavorable:

Requirements & Calculations

Final Answer

1.

Direct-material price variance.

(9.10-9.60)*9000 = 4500 U

2.

Direct-material quantity variance.

(3300*2-7000)*9.1 = 3640 U

3.

Direct-labor rate variance.

(22.50-24)*13000 = 19500 U

4.

Direct-labor efficiency variance.

(3300*4-13000)*22.50 = 4500 F